OpenStudy (anonymous):

Maria rented a coat at $285 for 3 days. If she rents the same coat for 6 days, she has to pay a total rent of $510. Write an equation in the standard form to represent the total rent (y) that Maria has to pay for renting the coat for x days.

OpenStudy (anonymous):

standard form is Ax+By=C

OpenStudy (chris911):

Let (3,285) and (6,510) reprsent the coordinate points, where x = number of days coat is rented for; and y = price of rental Given these 2 points, we could write the equation of the straight line. We need the slope. Slope = (510-285)/(6-3) = 225/3 = 75 So the slope is 75. So the equation of the line is y - 285 = 75(x - 3) y - 285 = 75x - 225 y = 75x + 60

OpenStudy (chris911):

y = a +bx a is fixed cost, while b is variable cost so equation 1 & 2 are as follows, 285 = a +3b 510 = a + 6b

OpenStudy (anonymous):

do you know how to write it in standard form?

OpenStudy (anonymous):

standard form is Ax+Bx=C

OpenStudy (chris911):

check these websites to help u sorry im a bit busy http://openstudy.com/study#/updates/5273395be4b077aaee84bc89 and http://openstudy.com/study#/updates/5294f135e4b01081cd1ea500